# Quantitative periodic homogenization of parabolic equations with large drift and potential

### v2

@@ Abstract
1 Introduction
@@
$\left\{\begin{aligned} \frac{\partial u_{\varepsilon}}{\partial t}-\operatorname{div}\left(A\left(\frac{x}{\varepsilon}\right)\nabla u_{\varepsilon}\right)+\frac{1}{\varepsilon}b\left(\frac{x}{\varepsilon}\right)\cdot\nabla u_{\varepsilon}+\frac{1}{\varepsilon^{2}}c\left(\frac{x}{\varepsilon}\right)u_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ u_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ u_{\varepsilon}(0,x)&=u_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (1.1)
@@
$u_{\varepsilon}(t,x)=e^{\frac{-\lambda t}{\varepsilon^{2}}}v_{\varepsilon}(t,x)\psi\left(\frac{x}{\varepsilon}\right)$ (1.2)
where $(\lambda,\psi)$ are an eigenpair associated with a certain exponential $Y$-periodic eigenvalue problem. The use of exponential $Y$-periodic eigenvalue problems in homogenization can be traced back to [7, page 383]. Allaire and Orive in [4, Lemma 2.1] (see also [8, Lemma 3], where Capdeboscq treats operators with no lower order terms) show the existence of a unique $\theta\in\mathbb{R}^{d}$ such that the eigenvalue problems (with $A^{\ast}$ denoting the adjoint of $A$):
$\left\{\begin{aligned} -\operatorname{div}_{y}({A}(y)\nabla\psi(y))+b(y)\cdot\nabla_{y}\psi(y)+c(y)\psi(y)=\lambda\psi(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi(y)e^{-2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$. (1.3)
@@
$\left\{\begin{aligned} -\operatorname{div}_{y}({A}^{\ast}(y)\nabla\psi^{\ast}(y))-\operatorname{div}_{y}(b(y)\psi^{\ast}(y))+c(y)\psi^{\ast}(y)=\lambda^{\ast}\psi^{\ast}(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi^{\ast}(y)e^{2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$. (1.4)
@@
    $\min_{Y}\{\psi,\psi^{\ast}\}\geq a>0\qquad\mbox{ and }\qquad\int_{Y}\psi(y)\psi^{\ast}(y)\,{\rm d}y=1.$ (1.5)
@@
    $\beta(y):=\psi(y)\psi^{\ast}(y)b(y)+\psi(y)A^{\ast}(y)\nabla_{y}\psi^{\ast}(y)-\psi^{\ast}(y)A(y)\nabla_{y}\psi(y)\quad\mbox{ for }y\in Y,$ (1.6)
@@
$\left\{\begin{aligned} \sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(\alpha\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)+\frac{1}{\varepsilon}\beta\left(\frac{x}{\varepsilon}\right)\cdot\nabla v_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{\varepsilon}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$. (1.7)
@@
$\sigma(y):=\psi(y)\psi^{\ast}(y)\quad\mbox{ and }\quad\alpha(y):=\sigma(y)A(y)\quad\mbox{ for a.e. }y\in Y.$ (1.8)
@@
$u_{\varepsilon}(0,x)=u^{0}_{\varepsilon}(x)=u_{0}(x)\psi\left(\frac{x}{\varepsilon}\right)$ (1.9)
This is sometimes referred to as the well prepared initial datum. Similar well preparedness assumptions on the data have been made in the literature [3, 4]. This choice results in strong convergence of the factor $v_{\varepsilon}$ to its homogenized limit. If one were to keep the data $u_{\varepsilon}(0,x)$ independent of the parameter $\varepsilon$, then the factorization (1.2) forces the initial datum for $v_{\varepsilon}$ to be
$v_{\varepsilon}(0,x)=\frac{u_{0}(x)}{\psi\left(\frac{x}{\varepsilon}\right)}.$ (1.10)
@@
All of this can be argued using the notion of two-scale convergence [1], for example. In this case, the convergence will be weak and hence one cannot talk about rates.
@@
$\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\sqrt{\varepsilon}\left\{\left\|\partial_{t}v_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|v_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|u_{0}\right\|_{L^{2}(\Omega)}\right\}.$ (1.11)
@@
Before we comment further on the proof of Theorem 1.2, we return to the equation (1.7). As $\beta$ (defined by (1.6)) is periodic, divergence-free and is of zero average, there exists a periodic skew-symmetric matrix-valued function $y\mapsto B(y)$ such that (see [8, Lemma 1, p. 809])
@@
$\sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(M\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)=0\qquad\mbox{ in }(0,T)\times\Omega.$ (1.12)
@@
where the ellipticity constant $\mathfrak{b}$ depends on the ellipticity constant $\mu$ of $A$ and the lower bound $a$ in (1.5). Using classical homogenization arguments (see [10, Theorem 11.4, p. 211]), we can show that
$\lim_{\varepsilon\to 0}\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}=0,$ (1.13)
@@
$\left\{\begin{aligned} \frac{\partial v_{0}}{\partial t}-\operatorname{div}\left(M_{h}\nabla v_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{0}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$. (1.14)
@@
$\left\{\begin{aligned} \operatorname{div}_{y}\left(M(y)\nabla_{y}\left(\chi_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\chi_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$. (1.15)
@@
$\left\|v_{\varepsilon}\right\|_{\mathrm{L}^{\infty}([0,T];\mathrm{L}^{2}(\Omega))}\leq C\left\|u_{0}\right\|_{\mathrm{L}^{2}(\Omega)}.$ (1.16)
@@
$\left\|u_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}=\int_{0}^{T}\int_{\Omega}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\left|\psi\left(\frac{x}{\varepsilon}\right)\right|^{2}\left|v_{\varepsilon}(t,x)\right|^{2}\,{\rm d}x\,{\rm d}t\leq C\int_{0}^{T}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\,{\rm d}t,$ (1.17)
@@
$\left\|u_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\leq C\varepsilon.$ (1.18)
As mentioned in Remark 1.1, in the absence of the well-prepared initial datum (1.9), the initial datum for $v_{\varepsilon}$ in (1.7) takes the form (1.10). Note that the uniform estimate (1.16) for $v_{\varepsilon}$ holds true even in this case. This implies that the above conditional (i.e. under the assumption of $\lambda>0$) decay rate (in $\varepsilon$) for the solution $u_{\varepsilon}$ to (1.1) remains true for general $\mathrm{L}^{2}$ initial data (i.e. no need for well-preparedness in the sense of (1.9)). More comments on the aforementioned conditional decay rate for $u_{\varepsilon}$ is relegated to our concluding remarks (see Section 4).
@@
Factorizations comparable to (1.2) have appeared before [11, 5, 4]. We would like to emphasise here that [11] treats the setting where both the $\varepsilon$-problem and the corresponding homogenized problem are posed in the full space, i.e. the spatial domain is the whole of $\mathbb{R}^{d}$. The authors of [5] too are working in a perforated medium built out of the whole of $\mathbb{R}^{d}$. This is inherently due to the effective non-zero drift present in the model. In [4], the authors pose the problem in a bounded spatial domain $\Omega$. However, they make a change of unknown which takes the domain to the whole space $\mathbb{R}^{d}$ and incidentally the homogenized equation obtained in [4, Theorem 3.1] is posed in $\mathbb{R}^{d}\times[0,T)$. To the best of our knowledge, this present work is the first time where parabolic problems with large drift are treated with both the $\varepsilon$-problem and the homogenized problems are posed in $\Omega_{T}$ with $\Omega$ being a bounded domain in $\mathbb{R}^{d}$.
@@
The results of this paper also yields quantitative rates for the periodic homogenization of parabolic equations with nondivergence form elliptic part and with large lower order terms. We draw inspiration from the work of Wenjia Jing and Yiping Zhang [13] which derived quantitative rates for the periodic homogenization of nondivergence form elliptic equations with large drift terms. The key idea in [13] is to use the invariant measure associated with the underlying diffusion process. This is a well-known procedure (see [6, 7], also see [9, section 2.2]) to transform nondivergence form equations into divergence form. This present work deviates from the earlier works by not appealing to this invariant measure. We demonstrate that via exponential eigenvalue problems akin to (1.3)-(1.4), one can transform the parabolic equation with terms in nondivergence form into a parabolic equation with terms in divergence form. The initial-boundary value problem, for the unknown $\rho_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, that we examine is the following:
$\left\{\begin{aligned} \frac{\partial\rho_{\varepsilon}}{\partial t}-\mathcal{K}\left(\frac{x}{\varepsilon}\right):\nabla^{2}\rho_{\varepsilon}+\frac{1}{\varepsilon}q\left(\frac{x}{\varepsilon}\right)\cdot\nabla\rho_{\varepsilon}+\frac{1}{\varepsilon^{2}}r\left(\frac{x}{\varepsilon}\right)\rho_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ \rho_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ \rho_{\varepsilon}(0,x)&=\rho_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$. (1.19)
@@
$\rho_{\varepsilon}(t,x)=e^{\frac{-\widehat{\lambda}t}{\varepsilon^{2}}}\varphi_{\varepsilon}(t,x)\widehat{\psi}\left(\frac{x}{\varepsilon}\right),$ (1.20)
@@
We finish this rather long introduction by detailing the structure of the rest of the paper. Section 2 quantifies the convergence (1.13) for general second order parabolic problems with the positive coefficient next to the time derivative term being merely measurable and bounded. The main result of section 2 is Theorem 2.1 which gives a rate of $\mathcal{O}(\varepsilon^{\frac{1}{4}})$. Section 3 improves this rate to $\mathcal{O}(\varepsilon^{\frac{1}{2}})$ with an additional assumption that the coefficient of the time derivative is regular (see Theorem 3.1). The manuscript concludes with an appendix which gathers some useful results on the parabolic smoothing operator.

### v3

@@ Abstract.
Key words and phrases:
Homogenization; non self-adjoint operators; convection-diffusion; periodic medium; Convergence rates.
1991 Mathematics Subject Classification:
Primary: 35B27, 47B28, 76R50, 35K57, 35P15; Secondary: 74Q10.
Kshitij Sinha
Department of Mathematics, Indian Institute of Technology Bombay, Mumbai, 400076, India
1. Introduction
@@
(1.1) $\left\{\begin{aligned} \frac{\partial u_{\varepsilon}}{\partial t}-\operatorname{div}\left(A\left(\frac{x}{\varepsilon}\right)\nabla u_{\varepsilon}\right)+\frac{1}{\varepsilon}b\left(\frac{x}{\varepsilon}\right)\cdot\nabla u_{\varepsilon}+\frac{1}{\varepsilon^{2}}c\left(\frac{x}{\varepsilon}\right)u_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ u_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ u_{\varepsilon}(0,x)&=u_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$.
@@
(1.2) $u_{\varepsilon}(t,x)=e^{\frac{-\lambda t}{\varepsilon^{2}}}v_{\varepsilon}(t,x)\psi\left(\frac{x}{\varepsilon}\right)$
where $(\lambda,\psi)$ are an eigenpair associated with a certain exponential $Y$-periodic eigenvalue problem. The use of exponential $Y$-periodic eigenvalue problems in homogenization can be traced back to [8, page 383]. Allaire and Orive in [4, Lemma 2.1] (see also [9, Lemma 3], where Capdeboscq treats operators with no zeroth order terms) show the existence of a unique $\theta\in\mathbb{R}^{d}$ such that the eigenvalue problems (with $A^{\ast}$ denoting the adjoint of $A$):
(1.3) $\left\{\begin{aligned} -\operatorname{div}_{y}({A}(y)\nabla\psi(y))+b(y)\cdot\nabla_{y}\psi(y)+c(y)\psi(y)=\lambda\psi(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi(y)e^{-2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$.
@@
(1.4) $\left\{\begin{aligned} -\operatorname{div}_{y}({A}^{\ast}(y)\nabla\psi^{\ast}(y))-\operatorname{div}_{y}(b(y)\psi^{\ast}(y))+c(y)\psi^{\ast}(y)=\lambda^{\ast}\psi^{\ast}(y)\qquad\mbox{ in }(0,1)^{d}\\ y\mapsto\psi^{\ast}(y)e^{2\pi\theta\cdot y}\quad\mbox{ is }Y\mbox{-periodic},\end{aligned}\right$.
@@
    (1.5) $\min_{Y}\{\psi,\psi^{\ast}\}\geq a>0\qquad\mbox{ and }\qquad\int_{Y}\psi(y)\psi^{\ast}(y)\,{\rm d}y=1.$
@@
    (1.6) $\beta(y):=\psi(y)\psi^{\ast}(y)b(y)+\psi(y)A^{\ast}(y)\nabla_{y}\psi^{\ast}(y)-\psi^{\ast}(y)A(y)\nabla_{y}\psi(y)\quad\mbox{ for }y\in Y,$
@@
(1.7) $\left\{\begin{aligned} \sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(\alpha\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)+\frac{1}{\varepsilon}\beta\left(\frac{x}{\varepsilon}\right)\cdot\nabla v_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{\varepsilon}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{\varepsilon}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$.
@@
(1.8) $\sigma(y):=\psi(y)\psi^{\ast}(y)\quad\mbox{ and }\quad\alpha(y):=\sigma(y)A(y)\quad\mbox{ for a.e. }y\in Y.$
@@
(1.9) $u_{\varepsilon}(0,x)=u^{0}_{\varepsilon}(x)=u_{0}(x)\psi\left(\frac{x}{\varepsilon}\right)$
This is sometimes referred to as the well prepared initial datum. Similar well preparedness assumptions on the data have been made in the literature [3, 4]. This choice results in strong convergence of the factor $v_{\varepsilon}$ to its homogenized limit. Note further that (1.9) is the same as
$u_{\varepsilon}(0,x)=u_{0}(x)e^{\frac{\theta\cdot x}{\varepsilon}}P\left(\frac{x}{\varepsilon}\right)$
for a $Y$-periodic function $P$. Well prepared initial data in this form have appeared in [4, Remark 3.2], [10, Theorem 4.4] and [3, Theorem 3.2]. If one were to keep the data $u_{\varepsilon}(0,x)$ independent of the parameter $\varepsilon$, then the factorization (1.2) forces the initial datum for $v_{\varepsilon}$ to be
(1.10) $v_{\varepsilon}(0,x)=\frac{u_{0}(x)}{\psi\left(\frac{x}{\varepsilon}\right)}.$
@@
All of this can be argued using the notion of two-scale convergence [1], for example. The authors of [5] have studied the case of $\varepsilon$-independent initial data with compact support. They consider a factorisation similar to (1.2) and find an asymptotic expression for the factor $v_{\varepsilon}$ using the representation formula involving the Green’s function. The analysis carried out in [5] is quite intricate and makes certain assumptions on the geometry of the support of the initial datum (see [5, page 305]).
@@
(1.11) $\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}$ $\leq C\sqrt{\varepsilon}\left\{\left\|\partial_{t}v_{0}\right\|_{L^{2}(0,T;W^{1,d}(\Omega))}+\left\|v_{0}\right\|_{\mathrm{L}^{2}(0,T;\mathrm{H}^{2}(\Omega))}+\left\|u_{0}\right\|_{L^{2}(\Omega)}\right\}.$
@@
Before we comment further on the proof of Theorem 1.2, we return to the equation (1.7). As $\beta$ (defined by (1.6)) is periodic, divergence-free and is of zero average, there exists a periodic skew-symmetric matrix-valued function $y\mapsto B(y)$ such that (see [9, Lemma 1, p. 809])
@@
(1.12) $\sigma\left(\frac{x}{\varepsilon}\right)\frac{\partial v_{\varepsilon}}{\partial t}-\operatorname{div}\left(M\left(\frac{x}{\varepsilon}\right)\nabla v_{\varepsilon}\right)=0\qquad\mbox{ in }(0,T)\times\Omega.$
@@
where the ellipticity constant $\mathfrak{b}$ depends on the ellipticity constant $\mu$ of $A$ and the lower bound $a$ in (1.5). Using classical homogenization arguments (see [12, Theorem 11.4, p. 211]), we can show that
(1.13) $\lim_{\varepsilon\to 0}\left\|v_{\varepsilon}-v_{0}\right\|_{\mathrm{L}^{2}(\Omega_{T})}=0,$
@@
(1.14) $\left\{\begin{aligned} \frac{\partial v_{0}}{\partial t}-\operatorname{div}\left(M_{h}\nabla v_{0}\right)&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ v_{0}&=0\qquad\mbox{ on }(0,T]\times\partial\Omega,\\ v_{0}(0,x)&=u_{0}(x)\qquad\mbox{ in }\Omega,\end{aligned}\right$.
@@
(1.15) $\left\{\begin{aligned} \operatorname{div}_{y}\left(M(y)\nabla_{y}\left(\chi_{j}(y)+y_{j}\right)\right)&=0\qquad\mbox{ in }Y,\\ \int_{Y}\chi_{j}(y)\,{\rm d}y&=0,\end{aligned}\right$.
@@
(1.16) $\left\|v_{\varepsilon}\right\|_{\mathrm{L}^{\infty}([0,T];\mathrm{L}^{2}(\Omega))}\leq C\left\|u_{0}\right\|_{\mathrm{L}^{2}(\Omega)}.$
@@
(1.17) $\left\|u_{\varepsilon}\right\|^{2}_{\mathrm{L}^{2}(\Omega_{T})}=\int_{0}^{T}\int_{\Omega}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\left|\psi\left(\frac{x}{\varepsilon}\right)\right|^{2}\left|v_{\varepsilon}(t,x)\right|^{2}\,{\rm d}x\,{\rm d}t\leq C\int_{0}^{T}e^{-\frac{2\lambda t}{\varepsilon^{2}}}\,{\rm d}t,$
@@
(1.18) $\left\|u_{\varepsilon}\right\|_{\mathrm{L}^{2}(\Omega_{T})}\leq C\varepsilon.$
As mentioned in Remark 1.1, in the absence of the well-prepared initial datum (1.9), the initial datum for $v_{\varepsilon}$ in (1.7) takes the form (1.10). Note that the uniform estimate (1.16) for $v_{\varepsilon}$ holds true even in this case. This implies that the above conditional (i.e. under the assumption of $\lambda>0$) decay rate (in $\varepsilon$) for the solution $u_{\varepsilon}$ to (1.1) remains true for general $\mathrm{L}^{2}$ initial data (i.e. no need for well-preparedness in the sense of (1.9)). More comments on the aforementioned conditional decay rate for $u_{\varepsilon}$ is relegated to our concluding remarks (see Section 4). We would like to mention that in [5], the authors have obtained a much better decay rate (albeit a conditional result) of order $\mathcal{O}(\varepsilon^{3}\varepsilon^{\frac{d-1}{2}})$ (see [5, Remark 5.2] for further details).
@@
Factorizations comparable to (1.2) have appeared before [13, 6, 4]. We would like to emphasise here that [13] treats the setting where both the $\varepsilon$-problem and the corresponding homogenized problem are posed in the full space, i.e. the spatial domain is the whole of $\mathbb{R}^{d}$. The authors of [6] too are working in a perforated medium built out of the whole of $\mathbb{R}^{d}$. This is inherently due to the effective non-zero drift present in the model. In [4], the authors pose the problem in a bounded spatial domain $\Omega$. However, they make a change of unknown which takes the domain to the whole space $\mathbb{R}^{d}$ and incidentally the homogenized equation obtained in [4, Theorem 3.1] is posed in $\mathbb{R}^{d}\times[0,T)$. In [5], however, the case where both the $\varepsilon$-problem and the homogenized problem are posed in $\Omega_{T}$ is treated. The present work is in similar spirit. Here, we try to quantify the convergence of $v_{\varepsilon}$ to the corresponding homogenized limit as opposed to getting an ansatz for $v_{\varepsilon}$ which is carried out in [5].
@@
The results of this paper also yields quantitative rates for the periodic homogenization of parabolic equations with nondivergence form elliptic part and with large lower order terms. We draw inspiration from the work of Wenjia Jing and Yiping Zhang [15] which derived quantitative rates for the periodic homogenization of nondivergence form elliptic equations with large drift terms. The key idea in [15] is to use the invariant measure associated with the underlying diffusion process. This is a well-known procedure (see [7, 8], also see [11, section 2.2]) to transform nondivergence form equations into divergence form. This present work deviates from the earlier works by not appealing to this invariant measure. We demonstrate that via exponential eigenvalue problems akin to (1.3)-(1.4), one can transform the parabolic equation with terms in nondivergence form into a parabolic equation with terms in divergence form. The initial-boundary value problem, for the unknown $\rho_{\varepsilon}:\overline{\Omega}\times[0,T)\to\mathbb{R}$, that we examine is the following:
(1.19) $\left\{\begin{aligned} \frac{\partial\rho_{\varepsilon}}{\partial t}-\mathcal{K}\left(\frac{x}{\varepsilon}\right):\nabla^{2}\rho_{\varepsilon}+\frac{1}{\varepsilon}q\left(\frac{x}{\varepsilon}\right)\cdot\nabla\rho_{\varepsilon}+\frac{1}{\varepsilon^{2}}r\left(\frac{x}{\varepsilon}\right)\rho_{\varepsilon}&=0\qquad\mbox{ in }(0,T)\times\Omega,\\ \rho_{\varepsilon}(t,x)&=0\qquad\mbox{ on }(0,T)\times\partial\Omega,\\ \rho_{\varepsilon}(0,x)&=\rho_{\varepsilon}^{0}(x)\qquad\mbox{ in }\Omega.\end{aligned}\right$.
@@
(1.20) $\rho_{\varepsilon}(t,x)=e^{\frac{-\widehat{\lambda}t}{\varepsilon^{2}}}\varphi_{\varepsilon}(t,x)\widehat{\psi}\left(\frac{x}{\varepsilon}\right),$
@@
We finish this rather long introduction by detailing the structure of the rest of the paper. Section 2 quantifies the convergence (1.13) for general second order parabolic problems with the positive coefficient next to the time derivative term being merely measurable and bounded. The main result of section 2 is Theorem 2.1 which gives a rate of $\mathcal{O}(\varepsilon^{\frac{1}{4}})$. Section 3 improves this rate to $\mathcal{O}(\varepsilon^{\frac{1}{2}})$ with an additional assumption that the coefficient of the time derivative is regular (see Theorem 3.1). Section 4 presents few scenarios where decay rate mentioned in Remark 1.4 holds true. The manuscript concludes with an appendix which gathers some useful results on the parabolic smoothing operator.